Curriculum Vitae:

Francesco Baldassarri

Born in Venezia, Italy, February 8, 1951. Lauréat de l’Alliance Française 1966. Laurea ”summa cum laude” in Mathematics from the University of Padova

3.4.1974, with a thesis on ”Interpretazione funzionale di certe iperalgebre e dei loro anelli di bivettori di Witt”, under the direction of Iacopo Barsotti.

Military service: June 1974 - August 1975. Married with Mariasilvia Mancini, 1 child Giovanni (18/10/94). Home adress: Piazza Capitaniato 12/A, 35139 PADOVA, ITALY. Work adress: Dipartimento di Matematica Pura Applicata, Università di

Padova, Via Belzoni 7, 35131 PADOVA, ITALY

Visiting Fellow at the Mathematics Department of Princeton University for two years, from 9.1.75 to 8.31.77.

Professore Incaricato of Geometry at the University of Ferrara in the year 1977/78 and Assistant Professor of Geometry at the University of Padova from 1.1.1979 to 11.1.1980.

Visiting Research Mathematician in Princeton, during the Fall term of 1978-79.

Professeur Associé at the University of Rennes during the Spring term of 1980, invited by P. Berthelot.

Full Professor of Geometry at the University of Trieste, from 11.1.1980. Visiting Fellow and Visiting Research Mathematician at Princeton Univer-

sity during the Fall term of 1981-82.

Full Professor of Geometry at the University of Padova from 11.1.1982 to the present date.

Dean of Graduate Studies in Pure Mathematics at the University of Padova for 5 years 1990-1994;

Professeur Associé at Paris 6, April 1983, invited by D. Bertrand. Invited speaker at the Taniguchi Symposium on Special Differential Equa-

tions held at Katata, August 1991; Professeur Associé at Paris 13, June 1993, invited by D. Barsky. ”Scientist in Charge” for the italian branch of the European Union Research

HCM Network “p-adic Methods in Arithmetic Algebraic Geometry” (Jan. 1, 1994-Dec, 31, 1997) and then of the TMR Network “Arithmetic Algebraic Geometry” (Sept. 1, 1996-);

Visiting Lecturer at Oklahoma State University at Stillwater, Sept. 1993. Professeur Associé at Paris 6, November 1995, invited by G. Christol. Fellow of the Japan Society for the Promotion of Science, April 1996. Visiting Lecturer at Oklahoma State University at Stillwater, Sept. 1997. Has been teaching courses of Geometry at various level, and specialistic

seminars on p-adic Analysis, differential equations, and cohomology of algebraic, complex-analytic and rigid-analytic varieties. In numerous visits abroad, has been invited by several research Istitutions such as I.A.S. and University in Princeton, I.H.P., Paris 6, Paris 13 in Paris, Luminy, Oberwolfach, RIMS of Kyoto, and the universities of: Münster, Bochum, Rennes, Strasbourg, Toulouse, Madrid, Barcelona, Minneapolis, Stillwater, Corvallis, Nijmegen, Groningen, Tokyo, Chiba, Kobe, Cortona, Milano, Bologna, Pisa, Trento.

Is Editor-in-Chief of the journal Rendiconti del Seminario Matematico dell’Università di Padova from January 1993 to the present date.

Is a member of the Istituto Veneto di Scienze Lettere ed Arti (Venice). Has organized the International Conference on p-adic Analysis, held in

Trento, Italy, 5.28.1989 - 6.3.1989, sponsored by the CIRM, with the participation of more than 80 specialists from all over the world.

Has organized the Instructional Conference on “p-adic methods in arithmetic algebraic geometry”, held at Levico Terme (Trento), Italy, 21 May - 3 June, 1995.

Has participated to the semester of Cohomologies p-adiques et Applications Arithmétiques”, organized by the Centre E. Borel, held at IHP, Paris, in Spring 1997.

Dec. 12,1997